### v1

# Underwater Image Enhancement via Dehazing and Color Restoration

Abstract
With the rapid development of marine engineering projects such as marine resource extraction and oceanic surveys, underwater visual imaging and analysis has become a critical technology. Unfortunately, due to the inevitable non-linear attenuation of light in underwater environments, underwater images and videos often suffer from low contrast, blurriness, and color degradation, which significantly complicate the subsequent research. Existing underwater image enhancement methods often treat the haze and color cast as a unified degradation process and disregard their independence and interdependence, which limits the performance improvement. Here, we propose a Vision Transformer (ViT)-based network (referred to as WaterFormer) to improve the underwater image quality. WaterFormer contains three major components: a dehazing block (DehazeFormer Block) to capture the self-correlated haze features and extract deep-level features, a Color Restoration Block (CRB) to capture self-correlated color cast features, and a Channel Fusion Block (CFB) to capture fusion features within the network. To ensure authenticity, a soft reconstruction layer based on the underwater imaging physics model is included. To improve the quality of the enhanced images, we introduce the Chromatic Consistency Loss and Sobel Color Loss to train the network. Comprehensive experimental results demonstrate that WaterFormer outperforms other state-of-the-art methods in enhancing underwater images.
Index Terms—  Degradation process, ViT model, Underwater image enhancement
1 Introduction
The ocean holds vast natural resources, so deep-sea exploration and development are crucial. Underwater imaging serves as a vital carrier of marine information, and the image quality greatly influences marine engineering. However, with its floating particles, light scattering, and selective light absorption, the deep ocean environment introduces noise, blur, and color distortion [1], which have detrimental effects on underwater detection, underwater visual perception, and other underwater tasks. Underwater image enhancement (UIE) aims to mitigate these issues, restore the true colors, enhance the details, and consequently increase the data reliability and usability. Therefore, research on UIE algorithms is necessary, urgent, and of great significance.
In the early days, specialized hardware [2], multiple different images [3] or special imaging methods [4] were used to enhance the quality of underwater images. Although these methods were effective in certain scenarios, their expensive hardware and complex conditions limited their widespread use. To achieve convenience and universality, increasingly many image processing and computer vision techniques are used to address the challenges of underwater imaging. These techniques can be broadly categorized into non-physical model-based methods, physical model-based methods, and data-driven methods. Non-physical model-based methods do not rely on underwater optical propagation models. They adjust the pixels of the image in the spatial or transform domain to improve the visual quality of the image [5, 6, 7]. Although non-physical model-based methods are convenient and easy to use, they have controversial fidelity due to the lack of physical priors. Moreover, under special conditions, these methods can cause more severe color distortion and artifacts. Physical model-based methods consider the process of light propagation underwater, make assumptions, and build corresponding mathematical models to enhance underwater images. Example physical model-based methods estimate the light scattering parameters [8], medium transmission parameters [9, 10]. Physical model-based methods are image restoration methods and offer higher fidelity. However, they can fail in complex underwater environments and challenging imaging conditions when the assumptions do not hold. In recent years, data-driven techniques have gained significant momentum, and more researchers have begun using deep learning to enhance underwater images [11, 12]. With their large parameter space and powerful learning capabilities, deep learning models often outperform traditional methods when ample datasets are available. However, data-driven methods have three challenges. First, most data-driven methods face a trade-off between performance and the number of parameters. Second, they separately consider features of the underwater environment or rely on a single degradation process to guide the network, which limits their UIE performance. Third, they often cannot guarantee the authenticity of the enhanced data.
Hence, we devised WaterFormer, which is a lightweight but high-performance network that considers the atomization and color cast as inherent primary degradation features to the underwater environment to restore images. Specifically, WaterFormer uses the DehazeFormer Block proposed in  [13] to achieve self-attention for the haze features and extract deep features. Using the channel-wise self-attention matrix transformation, our Color Restoration Block (CRB) performs self-attention on color cast features to help the network infer results with a more realistic color distribution. We also introduced the Channel Fusion Block (CFB), which synthesized the global information from different features and refined it to generate more effective fusion features. Considering the constraints of the underwater imaging physical model, we incorporated an underwater soft reconstruction layer at the end of the network to enhance the fidelity of the results. Furthermore, we introduced the Chromatic Consistency Loss and Sobel Color Loss to train the network in maintaining chromatic consistency, preserving fine color details, and enhancing the image quality and generalization across various datasets.
Overall, the main contributions of this work are as follows:
  • A novel end-to-end network to enhance underwater images is proposed. Our approach uses the DehazeFormer Block, successfully transforms haze features and improves the image deblurring capability. It introduces an underwater soft reconstruction layer to ensure the authenticity of the results.
  • We integrate the CRB and CFB with an encoder-decoder structure based on the global information between channels to generate a more realistic color distribution and accurately fuse features from different stages.
  • We propose the Chromatic Consistency Loss and Sobel Color Loss to improve the training process of the network and enhance its generalization ability.
2 Related Work
2.1 Traditional Methods
Traditional methods can be broadly categorized into methods based on non-physical models and those based on physical models. Non-physical-model-based methods do not involve modeling the underwater optical transmission process. Instead, they adjust the pixels of an image in the spatial domain or transform domain using the characteristic statistics and visual perceptual analysis. These methods include contrast enhancement [14], histogram equalization [15, 16], white balance [17], wavelet transform [18], fusion-based methods [19, 20], retinex-based methods [21, 22, 23], and others. For example, Hitam et al. [16] applied the mixture Contrast Limited Adaptive Histogram Equalization to underwater images in the RGB and HSV color spaces. Singh et al. [18] performed a discrete wavelet transform on images and estimated the color correction factors that were associated with specific color casts to adjust the input image pixels. Ancuti et al. [19] separately conducted color correction and contrast enhancement on underwater images and subsequently fused the two versions based on automatically calculated weights. Although most non-physical model-based methods are simple to implement and computationally efficient, they lack prior knowledge of the optical propagation, which can result in oversaturated or undersaturated color distributions.
Physical-model-based methods are commonly used in underwater image enhancement. They involve mathematically modeling the underwater degradation process, estimating the parameters of the degradation model, and ultimately recovering clear images. These methods address various aspects of underwater image degradation, such as estimating the parameters related to the light scattering [8] and medium transmission [9, 24]. For example, Hou et al. [8] used optical response functions, including Point Spread Function (PSF) and Modulation Transfer Function (MTF), and deconvolution techniques for image restoration. Ke et al. [10] considered the color, saturation, and detail information in images to construct scene depth maps and edge maps to estimate the medium transmission rates. Galdran et al. [25] adjusted the Dark Channel Prior (DCP) method by incorporating red channel priors to estimate the medium transmission map for underwater scenes. Wang et al. [24] introduced the adaptive attenuation-curve prior to estimate medium transmission parameters. Physical-model-based methods are based on prior assumptions and offer high fidelity in image restoration. However, their applicability is often limited to specific underwater degradation phenomena, so they have lower generality.
2.2 Data-driven Methods
The use of deep learning techniques in UIE tasks has been a major trend in recent years. These approaches mainly follow two paradigms: the encoding-decoding paradigm and the generative paradigm. On one hand, methods based on the encoding-decoding paradigm typically encode images into a latent space and continuously learn low-level features and deep representations of images to reconstruct the final results through a decoder. Encoding-decoding methods often use global residual connections or gate fusion mechanisms to facilitate network learning. Li et al. [26] introduced a gate fusion network called Water-Net, which enhanced images by fusing the original image with white-balanced, histogram-equalized, and gamma-corrected versions. Water-Net can adapt to different underwater environments and enhance the image. Xue et al. [27] analyzed underwater degradation factors from two perspectives: color distortions and veil effects. They developed the Multi-Branch Multi-Variable Network (MBANet) to restore underwater images. Li et al. [28] proposed an enhancement network called Ucolor, which operated in the HSV, RGB, and Lab color spaces to encode images and decodes image features with a medium transmission map. Therefore, Ucolor can learn and selectively enhance different visual representations of images.
On the other hand, Generative Adversarial Networks (GANs) are key components of the generative paradigm, and they mainly consist of a generator and a discriminator. The generator produces samples from random noise, and the discriminator distinguishes between real and generated samples. Through adversarial learning, they iteratively improve each other until the generator can generate realistic samples. Islam et al. [29] designed an adversarial network called FUnIE based on conditional generation, where the generator was constructed using the Unet [30] architecture, and the discriminator was a Markovian discriminator [31]. Cong et al. [32] proposed PUGAN, which guided the generator to produce more interpretable results by estimating medium transmission parameters. Liu et al. [33] split the network training into weak and strong supervision stages, which enhanced its performance while decreasing the dependence on paired training data.
Most studies on UIE tasks are based on Convolutional Neural Networks (CNNs) because they can effectively capture local features in images through convolutional and pooling layers. In recent years, the Visual Transformer (ViT) [34] was proposed, and it can better capture both local and global perspectives of images [35] and gain more similarity between features than CNNs. As a result, some researchers began using ViT to handle UIE tasks. For example, Peng et al. [36] used the ViT architecture to build a generative adversarial paradigm network called U-shape. Shen et al. [37] constructed a ViT network called UDAformer that included a spatial attention mechanism and a channel attention mechanism in an encoder-decoder paradigm.
Although many data-driven algorithms exhibit excellent generalization, they often grapple with a trade-off between parameter count and performance. These methods overlook the interconnected but independent nature of the haze and color distortion phenomena. Treating these features as a singular degradation process neglects the interplay between haze removal and color restoration. Furthermore, existing approaches lack tailored loss functions for underwater image enhancement (UIE), which diminishes the learning capacity of the network. Hence, we developed a lightweight, high-performance network for haze and color distortion features with two proposed loss functions to enhance the underwater image features and optimize the learning capability of the network.
3 Proposed Method
Fig. 1: Overall structure of WaterFormer, which includes the DehazeFormer Block as the core component for feature extraction and transformation in the network. We enhanced the ability of the network to aggregate local information by incorporating the FReLU activation function in the MLP structure. Additionally, our proposed CRB and CFB contribute to enhancing the color features of underwater images and integrating features from different branches. The entire WaterFormer is a three-stage improved Unet network, where the Down-Sample and Up-Sample stages are implemented using convolution and pixelshuffle, respectively.
3.1 Overall framework and hierarchical structure
The proposed WaterFormer adopts the U-shaped[30] encoder-decoder architecture as the backbone. In the encoder part, the network uses the DehazeFormer Block structure[13] with self-correlation to refine the haze features and deep features. The Color Restoration Block (CRB) was designed to enhance the color characteristics of the feature maps. In the decoder part, in addition to refined features, we used the CFB for skip connections to efficiently fuse different features at different stages. Finally, to improve the realism of the estimated image, we incorporated an underwater soft reconstruction layer into WaterFormer at the end, which innovated based on the soft reconstruction layer in [13] and replaced the original global residual learning. Specifically, the underwater image formation process in [38] is as follows:
${{U}_{\lambda}}(x)={{I}_{\lambda}}(x)\cdot{{T}_{\lambda}}(x)+{{A}_{\lambda}}\cdot(1-{{T}_{\lambda}}(x))$ (1)
where ${{U}_{\lambda}}(x)$ is the underwater image captured by the camera; ${{I}_{\lambda}}(x)$ is the clear latent image; ${{T}_{\lambda}}(x)$ is the transmission map; ${A}_{\lambda}$ is the global background light; ${\lambda}$ is the wavelength of light in the R, G, B channels. According to Eq. 1, the recovery process of the underwater image can be described as:
${{I}_{\lambda}}(x)={{U}_{\lambda}}(x)\cdot(\frac{1}{{{T}_{\lambda}}(x)}-1)+{{A}_{\lambda}}\cdot(1-\frac{1}{{{T}_{\lambda}}(x)})+{{U}_{\lambda}}(x)$ (2)
We set $K_{\lambda}(x)=\frac{1}{{{T}_{\lambda}}(x)}-1$ and $B_{\lambda}(x)={{A}_{\lambda}}\cdot(\frac{1}{{{T}_{\lambda}}(x)}-1)$ to obtain the simplified Eq. 3:
${{I}_{\lambda}}(x)={{U}_{\lambda}}(x)\cdot{{K}_{\lambda}}(x)-{{B}_{\lambda}}(x)+{{U}_{\lambda}}(x)$ (3)
This equation can be introduced as a prior into our network. Specifically, we drove the network to predict the variable $O\in{{\mathbb{R}}^{h\times w\times 6}}$ and decomposed the variable into $K_{\lambda}(x)\in{{\mathbb{R}}^{h\times w\times 3}}$ and $B_{\lambda}(x)\in{{\mathbb{R}}^{h\times w\times 3}}$. Eq. 3 is referenced to reconstruct the final ground truth estimation image.
To reduce the complexity and computational cost of the network, we removed a significant number of convolutional layers from the original DehazeFormer Block structure and only retained the convolution operations related to the self-attention and multilayer perceptron (MLP). To compensate for the limited utilization of local information due to the decrease in number of convolutional layers, we replaced the activation function in the MLP with a more locally enhanced FReLU [39] and designed WaterFormer as a three-stage Unet-based network. If there are more complex underwater environments and less demand for network speed performance, the number of encoding-decoding stages can be increased to enhance the network performance.
3.1.1 Color Restoration Block (CRB)
Due to the absorption and scattering of light in water, underwater images tend to appear bluish or greenish compared to land images, which is a channel-level per-pixel interference. Therefore, leveraging the global information among the channels for color balancing is an important task in underwater image enhancement. Inspired by the matrix transformations in  [40, 41], we designed the CRB to perform matrix transformations on the pixel-wise information among input feature channels using a cross-channel cross-covariance. This process generates an attention map that implicitly encodes the global context, facilitates interactions within the channels and promotes the color enhancement. The computational complexity of the channel-level global self-attention is linear, so the CRB can exhibit faster computational performance than spatial self-attention computations [42, 34, 35]. Fig. 2 illustrates the specific structure of the CRB.
Fig. 2: Overall structure of the CRB. Through channel-wise self-attention, WaterFormer can better self-transform and enhance color features in images.
We used the Rescale Layer Normalization (RLN) from [13] because it ensures consistency in the mean and variance between input and output feature maps. We believe that the RLN can help preserve the image brightness, restore colors, and reduce artifacts. For a given input $x\in{{\mathbb{R}}^{c\times h\times w}}$, we first applied the RLN and used the $1\times 1$ point-wise convolution and $3\times 3$ depth-wise convolution to map the RLN into the required $Q$, $K$, and $V$ values for self-attention computation. This process can be expressed by the following Eq. 4:
$Q,K,V=Flat(DConv(PConv(RLN(x))))$ (4)
Here, $PConv(\cdot)$ is a $1\times 1$ point-wise convolution; $DConv(\cdot)$ is a $3\times 3$ depth-wise convolution; $Flat(\cdot)$ is Flattening operation; $Q,K,V\in{{\mathbb{R}}^{c\times hw}}$.
During the computation of the global channel-wise self-attention, we performed a $3\times 3$ convolution operation on the $V$ values. This step helped improve the performance of the network by gathering information from the neighboring regions. The overall process of attention calculation can be described as follows:
${Atten}=(softmax(Q\cdot{{K}^{T}})/\gamma)\cdot Conv3(V)$ (5)
where $\gamma$ is the scaling coefficient for stable computation, and $Conv3(\cdot)$ is the convolution operation with a $3\times 3$ kernel size.
Finally, we restored the mean and variance (Affine) of $Atten$, added $Atten$ pixel-wise with the input $x$ and further refined the features through an MLP layer.
3.1.2 Channel Fusion Block (CFB)
Pixel-wise addition and concatenation are the most common strategies for multi-feature map fusion. However, these methods treat the information among different feature maps equally. We believe that due to the nonlinear absorption of light by water at different wavelengths, the information between channels in underwater image features is not equally important. Therefore, in our designed CFB, the information among different feature map channels is fused based on their importance. Fig. 3 shows the specific process of the CFB. For two given feature maps $x_{1}\in{{\mathbb{R}}^{c\times h\times w}}$ and $x_{2}\in{{\mathbb{R}}^{c\times h\times w}}$, we obtained the weights $\{{\alpha}_{1}$ and ${\alpha}_{2}\}$ using the global average pooling $GAP(\cdot)$, softmax function, and split operation. Then, we performed a weighted sum and refined the sum result using $PConv$, $DConv$, and residual connections. The more formal fusion process of the CFB is as follows:
$\{{{\alpha}_{1}},{{\alpha}_{2}}\}=split(softmax(GAP({{x}_{1}}),GAP({{x}_{2}})))$ (6)
${y}^{\prime}={{\alpha}_{1}}\cdot{{x}_{1}}+{{\alpha}_{2}}\cdot{{x}_{2}}$
$y={y}^{\prime}+DConv(PConv({y}^{\prime}))$
where $y\in{{\mathbb{R}}^{c\times h\times w}}$ is the fused generated feature map.
Fig. 3: Overall structure of the CFB.
3.2 Loss Function
The overall loss function of our WaterFormer contains three terms: the Reconstruction Loss ${{L}_{\ell_{1}}}$, Chromatic Consistency Loss ${{L}_{chroma}}$ and Sobel Color Loss ${{L}_{Sobel}}$.
Reconstruction Loss. For the given normalized real image ${{I}_{gt}}\in{{\mathbb{R}}^{h\times w\times c}}$ and network predicted image ${{I}_{pred}}\in{{\mathbb{R}}^{h\times w\times c}}$, ${{L}_{\ell_{1}}}$ is calculated as follows:
${{L}_{{{\ell}_{1}}}}=\frac{1}{h\times w\times c}\cdot\sum\limits_{k=1}^{c}{\sum\limits_{n=1}^{h}{\sum\limits_{m=1}^{w}{\left|{{I}_{gt}}(k,n,m)-{{I}_{pred}}(k,n,m)\right|}}}$ (7)
Chromatic Consistency Loss. To help the network learn the color differences between ${I}_{gt}$ and ${I}_{pred}$ while maintaining consistent chromaticity, we designed ${{L}_{chroma}}$ based on the chromatic similarity term proposed in [43]. First, we transformed ${I}_{gt}$ and ${I}_{pred}$ from the $RGB$ space to the $YIQ$ space using a matrix transformation:
$\left[\begin{matrix}Y\\ I\\ Q\\ \end{matrix}\right]=\left[\begin{matrix}0.299&0.587&0.114\\ 0.596&-0.274&-0.322\\ 0.211&-0.523&0.312\\ \end{matrix}\right]\left[\begin{matrix}R\\ G\\ B\\ \end{matrix}\right]$ (8)
The $I$ and $Q$ chromaticity channels of image ${I}_{gt}$ (${I}_{pred}$) are represented by ${I}_{gt}^{I}$ (${I}_{pred}^{I}$) and ${I}_{gt}^{Q}$ (${I}_{pred}^{Q}$), respectively. The chromatic consistency values ${{S}_{chroma}^{I}}$ and ${{S}_{chroma}^{Q}}$ of channels $I$ and $Q$ are calculated as follows:
${{S}_{chroma}^{I}}=\frac{\sigma(I_{gt}^{I},I_{pred}^{I})+{{c}_{1}}}{{{\sigma}^{2}}(I_{gt}^{I})+{{\sigma}^{2}}(I_{pred}^{I})+{{c}_{1}}}$ (9)
${{S}_{chroma}^{Q}}=\frac{\sigma(I_{gt}^{Q},I_{pred}^{Q})+{{c}_{2}}}{{{\sigma}^{2}}(I_{gt}^{Q})+{{\sigma}^{2}}(I_{pred}^{Q})+{{c}_{2}}}$
where $S_{chroma}^{I},S_{chroma}^{Q}\in\left(0,1\right]$, and the values closer to 1 indicate higher chromatic consistency between the two images. ${{\sigma}^{2}(\cdot)}$ is the variance calculation; ${{\sigma}(\cdot,\cdot)}$ is the covariance calculation; $c_{1}$ and $c_{2}$ are normalization constants to harmonize the chromatic consistency, which we set to 0.001 in our experiment. We defined $L_{chroma}$ as follows:
${L_{chroma}}=1-{S_{chroma}^{I}}\cdot{S_{chroma}^{Q}}$ (10)
In the experiment, to enhance the chromatic consistency of local regions, we set a sliding window with a size of $15\times 15$ and a stride of 1 to calculate the chromatic consistency loss for each local region. Finally, we obtained the total chromatic consistency loss by taking the mean of all local losses.
Sobel Color Loss. Hazing typically results in blurriness in underwater images, which manifests as the loss of edge details and a decrease in contrast. To overcome this challenge, we effectively preserved the fine color details between ${{I}_{gt}}$ and ${{I}_{pred}}$ by constraining the contour components of different color channels. Specifically, we defined the Sobel color loss in Eq. 11:
$L_{Sobel}^{x}=\frac{1}{3}\cdot\sum\limits_{\lambda}^{\{R,G,B\}}{{{L}_{{{\ell}_{1}}}}(Sobe{{l}_{x}}(I_{gt}^{\lambda}),Sobe{{l}_{x}}(I_{pred}^{\lambda}))}$ (11)
$L_{Sobel}^{y}=\frac{1}{3}\cdot\sum\limits_{\lambda}^{\{R,G,B\}}{{{L}_{{{\ell}_{1}}}}(Sobe{{l}_{y}}(I_{gt}^{\lambda}),Sobe{{l}_{y}}(I_{pred}^{\lambda}))}$
${{L}_{Sobel}}=L_{Sobel}^{x}+L_{Sobel}^{y}$
where $Sobe{{l}_{x}}(\cdot)$ and $Sobe{{l}_{y}}(\cdot)$ are the convolution operations of $Sobel$ operator in the $x$ and $y$ directions, respectively. The final loss function is defined as:
$L={{\lambda}_{1}}\cdot{{L}_{{{\ell}_{1}}}}+{{\lambda}_{2}}\cdot{{L}_{chroma}}+{{\lambda}_{3}}\cdot{{L}_{Sobel}}$ (12)
Here, ${\lambda}_{1}$, ${\lambda}_{2}$, and ${\lambda}_{3}$ are constants that balance the proportion of each loss term, which were set as 3, 1, and 3, respectively, with numerous experiments.
4 Experiments
In this section, we present the details and settings of the experiments. Then, we evaluate the performance of the proposed model and compare it with other representative methods. Finally, we conduct a series of ablation experiments to demonstrate the effectiveness of the proposed components.
4.1 Implementation Details
Our experiments were conducted on both synthetic underwater datasets [44] and real underwater dataset UIEBD [26]. In the synthetic underwater dataset, the reference images followed real color distributions, whereas the synthetic images simulated 10 different underwater degradation environments (I, IA, IB, II, and III simulated open sea underwater environments; 1, 3, 5, 7, and 9 simulated nearshore underwater environments). Each category of synthetic images had 1449 images, which were paired with reference images. We randomly selected 300 paired images from each category as the training set. The UIEBD dataset comprised 890 pairs of real underwater-reference images and 60 real underwater images without pairs. The reference images were obtained through enhancement using various traditional methods. We randomly partitioned 800 paired images as the training set. All images were resized to $256\times 256$ for training, and we augmented the dataset by applying random flips and rotations.
The experiments were conducted on tow NVIDIA GPUs (GeForce GTX 2080Ti) using the PyTorch 1.7.1 framework on an Ubuntu 16.04 system. We trained the network using the Adam optimizer with an initial learning rate of 0.001 and decayed it by a factor of $\frac{1}{2}$ every 50 epochs. We set the number of training iterations to 300 epochs and the batch size to 4.
4.2 Experiment settings
Benchmarks. In the synthetic underwater dataset, we randomly selected 30 additional paired images from each category of synthetic data as the test set, which was denoted as Test-S300. In the UIEBD dataset, we selected the remaining 90 paired images as the test set, which was denoted as Test-U90. Test-S300 and Test-U90 served as our full-reference benchmarks for the experiments. To demonstrate the generalization performance of our proposed network, we randomly selected 60 and 16 unpaired real underwater images from the UIEBD dataset and SQUID dataset [45], respectively, as the no-reference datasets. These datasets are called Test-U60 and Test-SQ16, respectively.
Compared Methods. We compared the proposed WaterFormer to other state-of-the-art underwater image enhancement methods based on deep learning: UWCNN[44], Water-Net[26], Ucolor[28], FUnIE[29], U-shape[36], UDAformer[37], and LE-Net[46]. In addition, we compared WaterFormer with the following traditional methods: Fusion-based[19], Retinex-based[21], MLLE[14], UDCP[47], RED[25], and HFM[48].
Evaluation Metrics. For Test-S300 and Test-U90, we used SSIM, PSNR, and NRMSE as the full-reference evaluation metrics for our experiments. These metrics are widely used to assess the similarity between two images due to their convenience and accuracy. Higher values of SSIM and PSNR or lower values of NRMSE indicate greater similarity between two images. For Test-U60, we used UCIQEE [49] and UIQM [50] as the no-reference evaluation metrics. Higher values of UCIQE and UIQM indicate better visual perception of the underwater images by humans. However, UCIQE and UIQM may not accurately reflect the actual visual perception in certain cases.
Test-S300 Test-U90 Overhead
Methods SSIM↑ PSNR↑ NRMSE↓ SSIM↑ PSNR↑ NRMSE↓ #Param↓ MACs↓ Time↓
Fusion-based[19] 0.651 12.65 0.685 0.872 21.80 0.188 × × 0.073s
Retinex-based[21] 0.641 14.54 0.520 0.773 17.34 0.283 × × 0.208s
MLLE[14] 0.603 12.38 0.667 0.740 16.50 0.315 × × 0.038s
UDCP[47] 0.597 13.84 0.500 0.689 13.85 0.418 × × 1.694s
RED[25] 0.659 12.66 0.674 0.833 19.03 0.244 × × 0.020s
HFM[48] 0.676 15.00 0.377 0.835 18.32 0.293 × × 0.328s
UWCNN[44] 0.640 14.62 0.456 0.818 18.10 0.265 0.041M 2.649G 0.002s
Water-Net[26] 0.746 19.25 0.322 0.905 23.17 0.153 1.091M 71.42G 0.016s
FUnIE[29] 0.769 21.34 0.231 0.814 18.94 0.239 7.020M 10.24G 0.002s
Ucolor[28] 0.813 22.23 0.235 0.897 21.23 0.180 148.8M 404.8G 0.169s
U-shape[36] 0.793 21.89 0.213 0.806 20.28 0.208 22.82M 2.983G 0.025s
UDAformer[37] 0.798 22.39 0.226 0.921 23.38 0.147 9.594M 41.59G 0.042s
LE-Net[46] 0.774 19.19 0.277 0.865 21.42 0.180 0.014M 0.690G 0.006s
WaterFormer 0.907 29.40 0.116 0.929 24.39 0.137 0.313M 7.749G 0.023s
Table 1: Quantitative comparison results for each method on the Test-S300 and Test-U90 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.3 Quantitative Comparisons
Full-reference Evaluation. Tab. 1 shows the quantitative comparison results on Test-S300 and Test-U90. The optimal results are highlighted in red, and the second-best results are highlighted in blue. WaterFormer achieved the best SSIM, PSNR, and NRMSE metrics on both full-reference test sets. Notably, for Test-S300, WaterFormer surpassed the second-best method by 9.4% in SSIM, 31.6% in PSNR, and shows a 9.7% lower NRMSE. This significant improvement indicated that previous methods were not able to effectively restore the true color distribution, detail distribution, and noise distribution of underwater images. Similarly, for Test-U90, where different methods were used to enhance the original images to obtain reference images, WaterFormer surpassed the second-best method by 0.8% in SSIM, 4.3% in PSNR, and exhibited a 1.0% lower NRMSE. This represents that our proposed network performs better in fitting underwater image enhancement compared with the previously best-performing methods.
Overhead. In Tab. 1, we included the inference speed of each method. Additionally, for deep learning methods, we measured the overall cost of the network using the total number of parameters (#Param) and multiply-accumulate operations (MACs). All measurements of the inference speed and network cost were conducted on images of size $256\times 256$. Compared to other methods, our proposed WaterFormer achieves optimal inference results while maintaining fast inference speed and relatively low network overhead.
Non-reference Evaluation. Tab. 2 shows the quantitative comparison results on Test-U60 and Test-SQ16. The best-performing and second-best results are highlighted in red and blue, respectively. Unfortunately, our proposed WaterFormer did not achieve the best or second-best results in the UIQM and UCIQE metrics on either no-reference datasets. As referenced  [28, 36] notes, while the UIQM and UCIQE metrics effectively evaluate the color, brightness, contrast, saturation, and sharpness of underwater images, they struggle to accurately measure color artifacts and shifts. In scenarios with more color artifacts and shifts, these metrics tend to yield higher scores, contradicting human perception. Our WaterFormer consistently avoids color artifacts and shifts, contributing to our lower UIQM and UCIQE scores. Hence, while we consider UIQM and UCIQE scores as indicative, they shouldn’t be the sole criteria for assessing the algorithm performance.
Test-U60 Test-SQ16
Methond UIQM↑ UCIQE↑ UIQM↑ UCIQE↑
Input 2.042 0.478 0.813 0.393
Fusion-based[19] 2.601 0.621 1.629 0.589
Retinex-based[21] 2.677 0.582 2.336 0.597
MLLE[14] 1.967 0.593 2.340 0.558
UDCP[47] 1.639 0.554 1.067 0.599
RED[25] 2.330 0.559 1.362 0.560
HFM[48] 2.298 0.585 2.273 0.607
UWCNN[44] 2.411 0.487 2.018 0.404
Water-Net[26] 2.694 0.598 2.562 0.567
FUnIE[29] 2.927 0.560 2.599 0.491
Ucolor[28] 2.695 0.564 2.383 0.508
U-shape[36] 2.812 0.528 2.184 0.493
UDAformer[37] 2.807 0.576 2.487 0.519
LE-Net[46] 2.549 0.561 2.297 0.512
WaterFormer 2.546 0.594 2.372 0.543
Table 2: Quantitative comparison results for each method on the Test-U60 and Test-SQ16 datasets. Red and blue colors indicate the $1^{st}$ and $2^{nd}$ ranks, respectively.
4.4 Qualitative Comparisons
Full-reference Evaluation. Fig. 4 and Fig. 5 show the qualitative comparison results on Test-S300 and Test-U90, respectively. For the Test-S300 dataset, the images enhanced by Fusion-based and RED methods had a reddish bias and failed to recover the true image details from the haze. The Retinex-based and UWCNN methods reduced the saturation of the enhanced images and even distorted the true color tones. Water-Net and Ucolor approximated the real colors of the synthesized images, but they still introduced significant artifacts. In contrast, WaterFormer did not generate artifacts and closely restored the brightness, color tones, and details in each region of the recovered images. Regarding the Test-U90 dataset, apart from Water-Net and WaterFormer, other methods failed to recover the desired colors, introduced artifacts, or could not effectively dehaze the images. Our proposed WaterFormer could better restore the colors and details and produced more accurate results than Water-Net.
Non-reference Evaluation. Fig. 6 and Fig. 7 show the qualitative comparison results on Test-U60 and Test-SQ16 , respectively. For the Test-U60 dataset, none of the methods satisfactorily performed. For example, the Fusion-based method generated artificial colors, the UDCP method produced low-brightness results, the UWCNN method shows minimal enhancement effect, the Water-Net method had limited effect on the shark image in the third row, the FUnIE method tended to produce color artifacts, and the U-shape method yielded low contrast results. Meanwhile, the images enhanced by WaterFormer achieved satisfying brightness, contrast, and color. Similarly, for the Test-SQ16 dataset, the Fusion-based, UDCP, UWCNN, and FUnIE methods failed to effectively restore the image colors. Water-Net and FUnIE introduced color shifts and artifacts, whereas the U-shape method had disappointing contrast and edge details in the enhanced images. In contrast, the images enhanced by WaterFormer restored as many colors and details as possible but exhibited minimal color shifts and artifacts.
[…]
Fig. 4: Qualitative Comparison Results on Test-S300. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
[…]
Fig. 5: Qualitative Comparison Results on Test-U90. (a) Input, (b) Fusion-based[19], (c) Retinex-based[21], (d) RED[25], (e) UWCNN[44], (f) Water-Net[26], (g) Ucolor[28], (h) WaterFormer, (i) GT.
[…]
Fig. 6: Qualitative Comparison Results on Test-U60. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
[…]
Fig. 7: Qualitative Comparison Results on Test-SQ16. (a) Input, (b) Fusion-based[19], (c) UDCP[47], (d) UWCNN[44], (e) Water-Net[26], (f) FUnIE[29], (g) U-shape[36], (h) WaterFormer.
[…]
Fig. 8: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-S300, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
[…]
Fig. 9: Rows 1-4 show the enhanced results, histogram display results, Canny edge detection results, and saliency object detection results on Test-U90, respectively. (a) Input, (b) MLLE[14], (c) HFM[48], (d) Ucolor[28], (e) UDAformer[37], (f) LE-Net[46], (g)WaterFormer, (h) GT.
4.5 Other Comparisons
To further demonstrate that our proposed method has better color correction and broader applications than other methods, qualitative analyses were conducted on Test-S300 and Test-U90. Each image underwent histogram display, Canny edge detection, and saliency object detection. The saliency object detection method was referenced from [51], and Fig. 8 and Fig. 9 show the corresponding results. For Test-S300 and Test-U90, with the exception of WaterFormer, all methods failed to adequately reproduce the color distribution, gradient distribution, and salient feature distribution of the true/reference images. For example, the UDAformer generated color artifacts on Test-S300, which incurred inaccurate edge detection results. Meanwhile, when applied to Test-U90, MLLE uniformly processed all channels, which introduced gradient noise and diminished the salient features of the image. In summary, WaterFormer exhibited outstanding performance in the UIE task and proves to be the most beneficial in related applications.
4.6 Ablation Study
The proposed WaterFormer incorporates three major components: the CRB, CFB, and loss of combining ${{L}_{chroma}}$ and ${{L}_{Sobel}}$. To demonstrate the effectiveness of each component, an ablation study was performed on the baseline framework with various combinations of these components. Tab. 3 shows the statistical results on the Test-S300 dataset, where the full model ($V_{5}$) outperformed other variants in terms of SSIM and PSNR scores.
Variants CRB CFB ${{L}_{chroma}}$ ${{L}_{Sobel}}$ SSIM↑ PSNR↑
Base w/o w/o w/o w/o 0.872 26.64
${{V}_{1}}$ w/o w/o w/o 0.876 27.09
${{V}_{2}}$ w/o w/o 0.877 27.26
${{V}_{3}}$ w/o 0.891 28.47
${{V}_{4}}$ w/o 0.901 28.56
${{V}_{5}}$ 0.907 29.40
Table 3: Quantitative analysis of the WaterFormer ablation experiments. The red font indicates the best performance.
Fig. 10 displays the visual results of various variants of WaterFormer. The Base model has basic UIE capabilities but retained significant shortcomings in addressing the color degradation and edge artifacts. The $V_{1}$ model, which incorporates the CRB module, could more delicately restore colors. However, due to insufficient feature fusion, the results of the $V_{1}$ model remained lacking. Therefore, by introducing the CFB module, the predictions of the $V_{2}$ model more closely aligned with the true color distribution. The addition of ${{L}_{chroma}}$ in the $V_{3}$ model further enhanced the consistency between inference results and chromaticity of the real images. However, $V_{1}$, $V_{2}$, and $V_{3}$ continued to exhibit noticeable edge artifacts, and this problem was addressed by the inclusion of ${{L}_{Sobel}}$, as demonstrated in the results of the $V_{4}$ inference. Finally, by combining all components, our WaterFormer ($V_{5}$) yielded the closest results to the real images.
[…]
Fig. 10: Qualitative analysis of the WaterFormer ablation experiments. (a) Input, (b) Base, (c) $V_{1}$, (d) $V_{2}$, (e) $V_{3}$, (f) $V_{4}$, (g) $V_{5}$, (h) GT.
In addition, we will demonstrate the effectiveness of other improvements on WaterFormer in Tab. 4, including replacing the ReLU with the FReLU [39] in the MLP structure and replacing the global residual reconstruction layer with the soft reconstruction layer that was proposed in [13] and optimized according to Eq. 3. References [13, 52] also proposed the SK Fusion structure to fuse features from different stages. Unfortunately, we found that the SK Fusion structure was not suitable for feature fusion in underwater images, possibly because SK Fusion does not effectively utilize the global information among different feature channels in underwater images. We present the ablation experiments on SK Fusion in Tab. 4.
Variants SSIM↑ PSNR↑
WaterFormer 0.907 29.40
FReLU [39]→ReLU 0.905 29.40
UW Soft Recon→Recon 0.904 29.00
UW Soft Recon→Soft Recon [13] 0.904 28.93
CFB→SK fuison [13, 52] 0.903 28.88
Table 4: Quantitative analysis of the ablation experiments on other improvements in WaterFormer. The red font indicates the best performance.
5 Conclusion
In this work, we introduced WaterFormer, a versatile method to enhance the quality of underwater images. To tackle challenges such as haziness and color shifts, WaterFormer leverages the Vision Transformer architecture and comprises three key components: the DehazeFormer Block, Color Restoration Block (CRB), and Channel Fusion Block (CFB). The DehazeFormer Block transforms hazy features and extract deep-level features. The CRB enhances colors through channel-level transpose matrices. The CFB facilitates feature fusion based on the importance of global information across different feature channels. We also incorporated the Chromatic Consistency Loss and Sobel Color Loss to ensure color fidelity and preserve fine details. The quantitative evaluations, qualitative evaluations, and ablation experiments show that our proposed network performs significantly better than various methods.
Nevertheless, this study has several limitations. First, our work requires a substantial amount of paired data for training, which poses a significant challenge for practical applications. Second, we did not consider the effects of noise such as environmental noise and camera noise during the image processing. Large-scale datasets that cover and fully consider complex types of image noise will be essential to train a more robust and generalizable model.
In our subsequent work, we will consider incorporating a denoising module into the network to reduce the impact of noise and enhance the robustness of the image correction results. In addition, we will use a contrastive learning strategy to reduce the network dependency on data and improve the generalization performance. The denoising process will be integrated into the end-to-end network as a module instead of a separate preprocessing step. We will also explore the processing or analysis based on enhanced results, such as the underwater image classification and measures of image representability. We believe that the analysis of restored underwater images can provide incremental assistance in future underwater visual-imaging tasks.